# Revisiting the Radial Metallicity Gradient-Age Relation in the Milky Way’s Thin and Thick Disks

### v1

@@ 4 Discussion
Figure 5: Comparison between our observed MGAR in the thin disk and literature results (solid lines for MGARs, dashed lines for stellar birth/ISM gradient evolutions).Upper panel: MGAR in the thin disk compared with reconstructed birth gradients from Ratcliffe et al. (2023) and Lu et al. (2024). The divergence between the MGAR and birth gradient at old ages may reflect the cumulative effect of churning. The gradient of the youngest stars aligns with the birth gradients ($\sim-0.07\,\mathrm{dex\,kpc^{-1}}$) since little migration could occur. Lower panel: MGAR compared with two results predicting steepening gradients: the gas-phase evolution from Sharda et al. (2021b) and the birth gradient from FIRE2 simulations (Graf et al., 2024), we also plot the MGAR from the FIRE2 simulations. See the text for further discussion.
@@ 4.1 Origin of the Flattening Gradient in the Thin Disk
As discussed in Section 1, radial migration—particularly churning—is a likely mechanism for redistributing stars after their formation, thereby contributing to the observed flattening of the MGAR (see Figure 3, upper panel). However, it remains uncertain whether the flattening primarily reflects stellar migration or intrinsic evolution in the ISM gradient over time.
@@
Several studies have explored this question. Within the inside-out formation framework, Minchev et al. (2018) reconstructed stellar birth radii and found a gradual flattening of the birth metallicity gradient over time. Similar flattening trends have been observed in external galaxies, where more massive systems tend to show shallower gradients. The analytic model of Sharda et al. (2021a) captures this mass dependence, predicting that gradients measured in $\mathrm{dex\,kpc^{-1}}$ flatten with increasing stellar mass, while gradients in $\mathrm{dex\,r_{e}^{-1}}$ follow a U-shape ($r_{e}$ is the effective radius). Interestingly, applying this model to MW–like galaxies yields opposite gradient evolutions depending on assumptions: a steepening ISM gradient over time, or a flattening gradient if the gas velocity dispersion is scaled by $(1+z)^{-1}$, to account for possible overestimates in star formation rates (Sharda et al., 2021b). These discrepancies highlight the need for more robust theoretical treatments.
@@
Recent FIRE2 simulations challenged the migration-driven scenario. Graf et al. (2025) found that the present-day MGAR closely follows the evolution of the stellar birth gradient (or ISM gradient then), suggesting minimal impact from radial migration. As shown in the simulations of Graf et al. (2024), gas-phase metals are efficiently mixed by turbulence and diffusion before star formation, suppressing spatial metallicity differences. As a result, prominent metallicity gradients appear only after the disk settles and turbulence decreases. This interpretation emphasizes gas-phase processes over stellar redistribution.
Figure 5 contrasts our observed MGAR with several of the previously mentioned studies (solid lines for MGARs, dashed lines for stellar birth/ISM gradient evolutions). In the upper panel, we compare our thin disk MGAR with the reconstructed birth gradients from Ratcliffe et al. (2023) and Lu et al. (2024). The MGAR steepens and the birth gradients flatten as stellar age decreases, and they converge near $0.07\,\mathrm{dex\,kpc^{-1}}$ for the youngest stars, where migration is negligible. Lu et al. (2024) proposed that the sharp steepening of birth gradient between 12 and 10 Gyr in Ratcliffe et al. (2023) could be attributed to the Gaia–Enceladus/Sausage merger (Helmi et al., 2018; Belokurov et al., 2018).
In the lower panel of Figure 5, we compare our MGAR in the thin disk to the gas-phase gradient evolution predicted by Sharda et al. (2021b) and to both the birth gradient and MGAR from FIRE2 simulations (Graf et al., 2024). Our observed MGAR agrees broadly with Sharda et al. (2021b) except for the upturn at younger ages ($\lesssim 3$ Gyr), which is likely due to enhanced dilution from late gas accretion. Within FIRE2, the similarity between the MGAR (green solid line) and the birth gradient (green dashed line) suggests minimal churning. Similar to other studies, our MGAR appears significantly steeper than that of FIRE2, possibly due to uncertainties in Galactic distance measurements or an earlier disk settling in the MW compared to FIRE2 galaxies or nearby MW analogs (Graf et al., 2024).
@@ 4.2 Origin of the Positive Gradient in the Thick Disk
[…]
Given the relatively small positive gradient we observe ($\sim 0.013\,\mathrm{dex\,kpc^{-1}}$), another possibility is that the thick disk initially had no radial gradient. This would align with the clumpy formation scenario of the thick disk, where giant clumps heat the galactic disk and violent perturbations erase chemodynamical relations during the formation of the thick disk (e.g., Inoue & Saitoh, 2014).
@@
Future studies could apply parameterized models such as the one used in Frankel et al. (2020) to improve the robustness of MGAR measurements. Such parameterized models and the multi-ring chemical evolution model in Johnson et al. (2021) may also help distinguish between different birth gradient evolution scenarios shown in Figure 5.
@@ 5 Conclusions
6 ACKNOWLEDGEMENTS
We thank the developers and maintainers of the following software libraries, which were used in this work: ArviZ (Kumar et al., 2019), Astropy (Astropy Collaboration et al., 2022), galpy (Bovy, 2015), matplotlib (Hunter, 2007), NumPy (Harris et al., 2020), pandas (pandas development team, 2020; Wes McKinney, 2010), PyMC (PyMC-Devs, 2025), seaborn (Waskom, 2021), and scipy (Virtanen et al., 2020).
The research presented here is partially supported by the National Natural Science Foundation of China under grant Nos. 12025302, 11773052, 11761131016; by China Manned Space Program with grant no. CMS-CSST-2025-A11; and by the “111” Project of the Ministry of Education of China under grant No. B20019. This work has made use of the Gravity Supercomputer at the Department of Astronomy, Shanghai Jiao Tong University.
7 Data Availability
@@ Appendix A Fitting MGAR Incorporating All Uncertainties
$\mathrm{[Fe/H]}_{\mathrm{obs}}\sim\widetilde{\mathcal{N}}(\mathrm{[Fe/H]}\mid\nabla\,R_{\mathrm{g}}+b,\Sigma^{2}),$ (A1)
@@
    $R_{\mathrm{g}}\sim\widetilde{\mathcal{N}}(R_{\mathrm{g}}\mid R_{\mathrm{g,obs}},\Delta R_{\mathrm{g}}^{2}),\quad R_{\mathrm{g}}\in[0,\,20]~\mathrm{kpc}.$ (A2)
@@
    $\Sigma^{2}\propto\sigma^{2}+\Delta\mathrm{[Fe/H]}^{2}.$ (A3)
  • Uncertainty in age: To account for $\Delta\tau$, we apply a weight based on the probability that the true age lies within the assigned age bin. Assuming a truncated normal distribution for the true age within $0,\,13]\,\mathrm{Gyr}$, the effective variance $\Sigma$ in Equation A1 becomes:
    $\Sigma^{2}=\dfrac{\sigma^{2}+\Delta\mathrm{[Fe/H]}^{2}}{\left[\displaystyle\int_{\mathrm{bin}}\widetilde{\mathcal{N}}(\tau\mid\tau_{\mathrm{obs}},\Delta\tau^{2})\,\mathrm{d}\tau\right]^{2}}.$ (A4)
@@
    $\nabla\sim\mathcal{N}(0,0.5^{2}),\quad b\sim\mathcal{N}(0,0.5^{2}),\quad\sigma\sim\mathrm{HalfNormal}(0,0.4^{2}).$
@@
Figure B2 explores the effect of purely geometric selections, dividing the sample into four $Z_{\mathrm{max}}$ bins: (0–0.25), (0.25–0.8), (0.8–2), and (2–6) kpc, with no chemical cut applied. The first and second columns of Figure B2 show that the observed MGAR flattening in the thin disk mainly comes from stars with low $Z_{\mathrm{max}}$, which may be due to the preferred migration for stars on the vertically coldest orbits (Solway et al., 2012; Vera-Ciro et al., 2014).

### v4

@@ 4.1 Origin of the flattening gradient in the thin disk
As discussed in Section 1, radial migration (particularly churning) is a likely mechanism for redistributing stars after their formation, thereby contributing to the observed flattening of the MGAR (see Figure 3, upper panel). However, it remains uncertain whether the flattening primarily reflects stellar migration or intrinsic evolution in the ISM gradient over time.
Figure 5: Comparison between our observed MGAR in the thin disk and literature results (solid lines for MGARs, dashed/dash-dotted lines for stellar birth/gas-phase gradient evolutions). Upper panel: MGAR in the thin disk compared with reconstructed birth gradients retrieved from Rbirth python package (calculated from age data from Anders et al. (2023) and the selection function corrected results from Khoperskov et al. (2025)). The divergence between the MGAR and birth gradient at old ages may reflect the cumulative effect of churning. The gradient of the youngest stars aligns with the birth gradients ($\sim-0.065\,\mathrm{dex\,kpc^{-1}}$), as little migration is expected to occur. Lower panel: MGAR compared with steepening gradients: the gas-phase metallicity gradient evolution model from Sharda et al. (2021b, dash-dotted lines and the grey shade. $\phi_{y}$ is the yield reduction factor of the model: larger $\phi_{y}$ for stronger ISM enrichment; smaller $\phi_{y}$ for metallicity losses through galactic winds.) and the birth gradient from FIRE2 simulations (Graf et al., 2024), we also plot the MGAR from the FIRE2 simulations. See the text for further discussion.
@@
Several studies have explored this question. Recent method for estimating $R_{\mathrm{birth}}$ introduced in Section 1 found a gradually flattening of the birth metallicity gradients with time (Ratcliffe et al., 2025a). Similar flattening trends have been observed in external galaxies, where more massive systems tend to show shallower gradients. The analytic model of Sharda et al. (2021a) captures this mass dependence, predicting that gas-phase metallicity gradients measured in $\mathrm{dex\,kpc^{-1}}$ flatten with increasing stellar mass, while gradients in $\mathrm{dex\,r_{e}^{-1}}$ follow a U-shape ($r_{e}$ is the effective radius). Interestingly, applying this model to MW–like galaxies yields opposite gradient evolutions depending on assumptions: a steepening ISM gradient over time in their Figure 10, or a flattening gradient in their Figure B1 if the gas velocity dispersion is reduced by $(1+z)$, to account for possible overestimates in star formation rates (Sharda et al., 2021b). These discrepancies highlight the need for more robust theoretical treatments.
@@
Recent FIRE2 simulations challenged the migration-driven scenario. Graf et al. (2025) found that the present-day MGAR closely follows the evolution of the stellar birth gradient (or ISM gradient then), suggesting minimal impact from radial migration. As shown in the simulations of Graf et al. (2024), gas-phase metals are efficiently mixed by turbulence and diffusion before star formation, suppressing spatial metallicity differences. As a result, prominent metallicity gradients emerge only after the disk reaches a settled state and turbulence diminishes. This interpretation emphasizes gas-phase processes over stellar redistribution.
Figure 5 contrasts our observed MGAR with several previously mentioned studies: solid lines for MGARs, dashed lines for stellar birth–gradient evolutions, and dot–dashed lines for gas–phase gradient evolutions.
4.1.1 Comparison with flattening birth gradient
In the upper panel of Figure 5, we compare our thin–disk MGAR with two reconstructed birth–gradient evolution trends retrieved from the Rbirth Python package, one based on the ages from Anders et al. (2023), another from the selection–function–corrected results of Khoperskov et al. (2025). The upper panel reveals a growing divergence toward older ages (i.e., larger lookback time): the reconstructed birth gradients steepen substantially, whereas our observed MGAR becomes progressively flatter. As a result, the difference between the two reaches $\geq 0.08\,\mathrm{dex\,kpc^{-1}}$ by $\sim 10\,\mathrm{Gyr}$.
This widening offset at old ages is most naturally interpreted as the cumulative effect of radial migration: stars born with a steep radial metallicity gradient can be redistributed over large radial ranges over several Gyrs, thereby flattening the present–day MGAR relative to the (much steeper) reconstructed birth gradients. Although the reconstructed curves also carry uncertainties and may be affected by systematics, the increased discrepancy toward older ages implies that migration is an important shaping mechanism for the oldest thin disk populations.
We further note that Lu et al. (2024) associated the sharp steepening in the reconstructed gradients around 10–12 Gyr with the Gaia–Enceladus/Sausage merger (Helmi et al., 2018; Belokurov et al., 2018). This event adds an additional complication when comparing reconstructed birth gradients with the present–day MGAR at large lookback times, but does not alter the qualitative conclusion that migration helps explain the observed divergence.
4.1.2 Comparison with steepening birth/gas gradient
In the lower panel of Figure 5, we compare the MGAR of our thin disk with both the MGAR and the birth gradients from FIRE2 simulations (Graf et al., 2024), and with the analytical evolution of the gas–phase metallicity gradient from Sharda et al. (2021b). In the analytical model, the parameter $\phi_{y}$ ($0\leq\phi_{y}\leq 1$) represents the yield reduction factor: larger $\phi_{y}$ values imply that most supernova-produced metals remain in the ISM, while smaller $\phi_{y}$ values correspond to stronger metallicity losses through galactic winds.
Our observed MGAR broadly follows the $\phi_{y}=1.0$ model for ages $\gtrsim 3$ Gyr, except for an upturn at younger ages ($\lesssim 3$ Gyr), which is likely caused by dilution from late gas accretion. This agreement at intermediate and old ages suggests that, in this regime, the observed MGAR largely tracks the long–term evolution of the ISM metallicity gradient, and that radial migration plays only a secondary role. However, because the Sharda et al. (2021b) model assumes cylindrical symmetry and omits the bar and spiral structures, which are key drivers of radial migration in the Milky Way (Sellwood and Binney, 2002; Roskar, 2010; Zhang et al., 2025; Marques et al., 2025), hence its predicted gas-phase gradients are likely steeper than in a fully realistic disk. Thus, the apparent agreement with the $\phi_{y}=1.0$ curve should be interpreted with caution.
For the FIRE2 simulations, both the MGAR and the birth gradients (green solid/dashed lines) lie between the $\phi_{y}=0.1$ and $\phi_{y}=0.3$ curves, substantially flatter than our observed MGAR, which is much closer to the $\phi_{y}=1.0$ curve. Thus, the vertical offset is plausibly explained by the stronger metal ejection in FIRE2, which consequently produces flatter ISM/birth gradients. Other possible contributors to the offset include uncertainties in distance measurements or the possibility that the real Milky Way experienced earlier disk settling than FIRE2 galaxies or nearby analogs (Graf et al., 2024).
Although the overall trends of MGAR and birth gradients in FIRE2 are similar (i.e., the agreement between the green solid line and green dashed line), which indicates limited radial migration within the simulations, the weak and short bars in FIRE2 (Ansar et al., 2025) suggest that the efficiency of churning may be underestimated relative to the real Milky Way. Therefore, the FIRE2 agreement between MGAR and the birth gradients reflects the underlying ISM evolution in the simulations rather than providing a direct constraint on the Milky Way’s migration efficiency.
@@ 4.2 Origin of the positive gradient in the thick disk
The thick disk likely formed in the early evolutionary stages of the MW, when the disk was gas-rich and dynamically turbulent. Under such conditions, it is plausible that a weakly positive radial metallicity gradient could emerge. Simulations from the FIRE2 suite suggest that strong stellar feedback, especially from early starbursts, can drive metal-enriched outflows into the outer disk, leading to such positive gradients (Sun et al., 2025). The accretion of pristine gas into the MW center during early epochs may also contribute to forming a positive metallicity gradient (Cresci et al., 2010; Curir et al., 2012). Recent JWST observations show positive metallicity gradient ($0.165\,\mathrm{dex\,kpc^{-1}}$) at z $\sim 3$ (Wang et al., 2022b). For even higher redshifts, Venturi et al. (2024) measured slightly positive gradients ($0.03\sim 0.14\,\mathrm{dex\,kpc^{-1}}$) in $6<z<8$, while Li et al. (2025) reported flattening negative gradients ($0.53\,\mathrm{dex\,kpc^{-1}}$ to $0.34\,\mathrm{dex\,kpc^{-1}}$) from z $\approx 7$ to z $\approx 6$ (their Figure 2) and suggested that the slightly positive gradients in Venturi et al. (2024) are resulted from merging systems.
Given the relatively small positive gradient we observe ($\sim 0.013\,\mathrm{dex\,kpc^{-1}}$), another possibility is that the thick disk initially had no radial gradient. This would align with the clumpy formation scenario of the thick disk, where giant clumps heat the galactic disk and violent perturbations erase chemodynamical relations during the formation of the thick disk (e.g., Inoue and Saitoh, 2014).
@@
Future studies could aim to correct for the selection function and apply parameterized models to improve the robustness of MGAR measurements. Such parameterized models and chemical evolution models mentioned in Section 1 may also help to distinguish between different birth gradient evolution scenarios shown in Figure 5.
@@ 5 Conclusions
Acknowledgements
We thank the anonymous referee for helpful and constructive suggestions that strengthened the quality of this work. We thank the developers and maintainers of the following software libraries, which were used in this work: ArviZ (Kumar et al., 2019), Astropy (Astropy Collaboration et al., 2022), galpy (Bovy, 2015), matplotlib (Hunter, 2007), NumPy (Harris et al., 2020), pandas (pandas development team, 2020; McKinney, 2010), PyMC (PyMC-Devs, 2025), seaborn (Waskom, 2021), and scipy (Virtanen et al., 2020).
The research presented here is partially supported by the National Natural Science Foundation of China under grant Nos. 12533004, 12025302, 11773052; by China Manned Space Program with grant no. CMS-CSST-2025-A11; and by the “111” Project of the Ministry of Education of China under grant No. B20019. This work has made use of the Gravity Supercomputer at the Department of Astronomy, Shanghai Jiao Tong University.
Data Availability
@@ Appendix A Fitting MGAR Incorporating All Uncertainties
$\mathrm{[Fe/H]}_{\mathrm{obs}}\sim\widetilde{\mathcal{N}}(\mathrm{[Fe/H]}\mid\nabla\,R_{\mathrm{g}}+b,\Sigma^{2}),$ (3)
@@
    $R_{\mathrm{g}}\sim\widetilde{\mathcal{N}}(R_{\mathrm{g}}\mid R_{\mathrm{g,obs}},\Delta R_{\mathrm{g}}^{2}),\quad R_{\mathrm{g}}\in[0,\,20]~\mathrm{kpc}.$ (4)
  2. 2.
@@
    $\Sigma^{2}\propto\sigma^{2}+\Delta\mathrm{[Fe/H]}^{2}.$ (5)
  3. 3.
    Uncertainty in age: To account for $\Delta\tau$, we apply a weight based on the probability that the true age lies within the assigned age bin. Assuming a truncated normal distribution for the true age within $0,\,13]\,\mathrm{Gyr}$, the effective variance $\Sigma$ in Equation 3 becomes:
    $\Sigma^{2}=\dfrac{\sigma^{2}+\Delta\mathrm{[Fe/H]}^{2}}{\left[\displaystyle\int_{\mathrm{bin}}\widetilde{\mathcal{N}}(\tau\mid\tau_{\mathrm{obs}},\Delta\tau^{2})\,\mathrm{d}\tau\right]^{2}}.$ (6)
  4. 4.
@@
    $\nabla\sim\mathcal{N}(0,0.5^{2}),\,b\sim\mathcal{N}(0,0.5^{2}),\,\sigma\sim\mathrm{HalfNormal}(0,0.4^{2}).$ (7)
@@
Figure B2 explores the effect of purely geometric selections, dividing the sample into four $Z_{\mathrm{max}}$ bins: (0–0.25), (0.25–0.8), (0.8–2), and (2–6) kpc, with no chemical cut applied. The first and second columns of Figure B2 show that the observed MGAR flattening in the thin disk mainly comes from stars with low $Z_{\mathrm{max}}$, which may be due to the preferred migration for stars on the vertically coldest orbits (Solway et al., 2012; Vera-Ciro et al., 2014). The slightly non-monotonic MGAR of stars with intermediate $Z_{\mathrm{max}}$ is also shown in V21 and Willett et al. (2023).